How would you solve this equation? -4(p+2)+8=2(p-1)-7p+15
Answer by mananth(16946)

-4(p+2)+8=2(p-1)-7p+15
-4p-8+8 = 2p-2-7p+15
-4p=-5p+13
add +5p
5p-4p=-5p+5p+13
p=13
...
m.ananth@hotmail.ca